OpenStudy (anonymous):

Amy and Elizabeth both work in sales. Each week Amy earns $18/h plus a $40 bonus, and Elizabeth earns $12/h plus a $100 bonus. How many hours must Amy and Elizabeth each work in a week in order to earn the same amount of money? For that many hours, how much money would each earn? They would each have to work hours, and they would earn $ .

OpenStudy (alexrobin13):

18t+40=12t+100 (Where t is hours) The left side is Amy, and the right represents Elizabeth.

OpenStudy (alexrobin13):

To solve for time isolate t. So 18t+40=12t+100 6t+40=100 6t=60 t=10

OpenStudy (anonymous):

thanks my dead hurts so i cant really think but i think i got it

OpenStudy (alexrobin13):

To find how much they earn put in 10 as t, for one of the equations. 18(10)+40=x 180+40=x x=220

OpenStudy (alexrobin13):

So They would each have to work 10 hours, and they would earn $220 .

OpenStudy (anonymous):

thank u so much

OpenStudy (alexrobin13):

no problem